# Towards understanding evolution of science through language model series

### v1

@@ III-B Pretraining AnnualBERT
The original BERT model is based on a bidirectional Transformer architecture and pretrained through two objectives: Masked Language Model (MLM) and Next Sentence Prediction (NSP). In MLM, $15\%$ of input tokens are masked, among which $80\%$ are replaced with the [MASK] token, $10\%$ are replaced with randomly selected tokens from the vocabulary, and the remaining $10\%$ stay unchanged. The objective is to predict masked tokens using cross-entropy loss. NSP predicts whether or not two masked sentences follow each other in the original text. Built on BERT, the RoBERTa model [41] improves the pretraining by eliminating the NSP objective, which has been found less useful, and by using dynamic masking, where masked tokens are changed for each epoch, as opposed to fixed masking used in BERT where masked tokens are fixed across epochs.
@@
With $\mathcal{M}_{2008}$, we perform, for each year $t>2008$, continual training of $\mathcal{M}_{t-1}$ on the corpus formed by papers published in $t$ for one epoch, resulting in $\mathcal{M}_{t}$. For instance, a continual training of $\mathcal{M}_{2008}$ on the corpus based on papers published in 2009 yields $\mathcal{M}_{2009}$, which is then used to form $\mathcal{M}_{2010}$ using papers in 2010, in general following:
$\mathcal{M}_{t}=\text{Continual-train}\left(\mathcal{M}_{t-1},\ \text{corpus}_{t}\right)\,.$ (1)
Through this approach, we allow each instance of our AnnualBERT to capture evolving language and terminology specific to that year’s publications, so that we can represent the knowledge of a specific year in a more sophisticated way. The corpora used for pretraining our AnnualBERT grow to 2.7 billion words by the time we extend the model to AnnualBERT2020.
IV Performance on downstream NLP tasks
IV-A Task and dataset descriptions
@@
  1. 1.
    The SciERC dataset [42] annotates entities, relations, and coreference clusters in 500 abstracts from 12 AI conference and workshop proceedings indexed in the Semantic Scholar Corpus. It includes a total of $8,089$ distinct named entities.
  2. 2.
    The ScienceExam dataset [43] comprises entities from science exam questions aimed at students from 3rd to 9th grade, totaling $133,000$ entities.
  3. 3.
    The ScienceIE dataset [44] provides $9,946$ entities appeared in 500 scientific paragraphs.
@@
  1. 1.
    The SciERC dataset [42] includes $4,716$ annotated sentences categorized for classification according to relations such as USED-FOR, FEATURE-OF, PART-OF, etc.
  2. 2.
    The ACL-ARC dataset [45] comprises $1,941$ citation instances from 186 papers in the ACL Anthology Reference Corpus, with each instance annotated by domain experts with categories such as Future, Background, Motivation, etc.
  3. 3.
    The SciCite dataset [46] encompasses $11,020$ citation instances from $6,627$ scholarly papers, with each citation classified into categories such as Background, Method, and Result Comparison.
@@
The last task is motivated by the increasing prevalence of interdisciplinary science [47] and its role in tackling important societal problems [48].
@@
TABLE II: Experimental results of NLP tasks using different (scientific) BERT variants. Reported are F1 scores.
Task Field Dataset BERT BioBERT SciBERT BioMedBERT AnnualBERT
CLS CS SciERC 85.96 87.18 87.21 86.96 87.60
CS ACL-ARC 76.22 75.97 80.32 79.53 79.88
Multi SciCite 85.20 86.63 86.33 86.33 85.77
Multi arXiv (major) 88.83 88.96 90.69 89.95 91.41
Multi arXiv (sub) 44.48 45.31 57.03 51.60 58.24
Cross-field identification arXiv 83.58 85.42 87.85 85.82 92.13
NER CS SciERC 59.69 63.61 62.55 61.84 60.96
Multi ScienceExam 80.53 82.88 82.13 80.69 77.69
Multi ScienceIE 33.56 34.86 33.97 34.40 33.95
REL Multi ArXiv 86.79 87.64 89.18 87.44 89.52
IV-B Experimental setup
We compare our AnnualBERT models with other scientific BERT models, including BioBERT [4], SciBERT [6], BioMedBERT [7], as well as the original BERT [49]. For our evaluations, we utilize the AnnualBERT model corresponding to the dataset’s year of release.
To ensure consistent comparisons across various BERT variants, we standardize the training configuration for each model. For the NER tasks, we adopt the NER model structure from SimpleTransformers, which incorporates a direct linear layer, and set the training parameters as follows: maximum input length of 512, training duration of 10 epochs, and batch size of 32. The best model parameters after each epoch are saved. For the CLS tasks, the vectors obtained using the [CLS] token pooling strategy are used as the input into a multilayer perceptron (MLP) for classification during the fine-tuning phase. The primary training parameters are set as follows: a training duration of 4 epochs, a batch size of 16, and weight decay at 0.01, and the best-performing model on the development set is saved every 100 steps. For the REL task, we use the dual tower structure [50]. We feed the abstracts of two arXiv papers into the identical BERT-like model and employ the [CLS] tag pooling strategy to obtain their respective feature vectors, $u$ and $v$, that effectively represent each abstract. We then concatenate $u$ and $v$ and their element-wise absolute difference $\left|u-v\right|$. This concatenated vector is then fed into a MLP layer for classification.
@@ IV-C Experimental results
Table II presents the experimental results. Firstly, our AnnualBERT does not perform well on the NER tasks, and interestingly, neither do the other two models pretrained from scratch (SciBERT and BioMedBERT). This might due to the tokenization strategy we use during the pretraining process. It may be also related to the fact that BioBERT, the best-performing model, is resulted from continual pretraining from the general Wikipedia and books corpus, which may have a better alignment with the benchmark datasets like the ScienceExam.
@@
On the other hand, our AnnualBERT has the best performance in the CLS task on the SciERC dataset and comparable performances to the other specialized BERT models on the ACL-ARC and SciCite datasets. In general, the performance differences among these models in the three tasks are rather small. However, significant performance gaps become evident for the three domain-specific tasks, where our AnnualBERT performs the best. Particularly, for the sub-category prediction task, which involves classifying papers into one of the 172 sub-fields based on their abstracts, our AnnualBERT significantly leads the other models. Notably, SciBERT also has a comparable performance in this task, likely due to its training corpus containing a substantial number of computer science papers. Finally, AnnualBERT has a dominant performance in the cross-field paper identification task, indicating its superior knowledge representation ability for interdisciplinary papers that stretch multiple domains, while the other models, despite fine-tuned, struggle with this task. Taken together, these results demonstrate the validity of our AnnualBERT models and the effectiveness of domain-specific pretraining for domain-specific tasks, in consistence with prior studies [7].
V Citation link prediction
@@
Another perspective to assessing capabilities of our AnnualBERT models in learning and encoding semantic information of papers is through link prediction task in citation network. The rationale behind using this task lies in the nature of scientific discourse, where citations are formed by not only topical alignment but also nuanced interplay of numerous other factors like ideas and methodologies that are shared between two papers. Therefore, the ability of a language model in encoding abstracts into feature vectors that are highly predictive of whether two papers are related through citations reflects the model’s proficiency in grasping those complex semantic relationships, which in turn are instrumental for generating meaningful insights into domain-specific research [51].
@@
As our setup requires a citation network between arXiv papers, we derive such a network from the Microsoft Academic Graph (MAG) dataset [52]. Doing so can avoid the otherwise extensive efforts of matching each referenced entry within the arXiv corpus, which may involve identifying the published version of a arXiv paper. To build our citation network, we first identify the corresponding MAG versions of each arXiv paper, by extracting arXiv IDs from paper URLs provided by MAG, and then use the citation relationships between MAG papers to construct the network. Through this method, we identify nearly 1.5 million arXiv papers and 14.8 million citation relationships between them.
@@
V-A Link prediction in static network
@@
V-A1 Methods
Link prediction in networks has been extensively studied in the past [53, 54]. Here we consider two families of approaches to predicting missing links: topological predictors and graph neural networks (GNN).
@@ V-A1 Methods
  4. 4.
    The Adamic/Adar (AA) index [53] defines the score as $s=\sum_{z\in\Gamma(u)\cap\Gamma(v)}\frac{1}{\log\left|\Gamma(z)\right|}$. It considers the importance of a common neighbor $z$, by giving more weight if $z$ has fewer connections, capturing the intuition that common neighbors in a less-populated field can be more significant in predicting a link between $u$ and $v$.
  5. 5.
    The resource allocation (RA) index [55] defines the score as $s=\sum_{z\in\Gamma(u)\cap\Gamma(v)}\frac{1}{\left|\Gamma(z)\right|}$. It measures the likelihood of a link between two nodes based on the principle of resource allocation, where nodes share resources through their common neighbors.
  6. 6.
@@
One limitation of topological predictors is that they are unable to exploit node features, which in our context are the abstract of a paper. The rapid development of GNN has made it possible to integrate graph topology and node features for downstream tasks like link prediction. Here we leverage GraphSAGE [56] to generate embeddings $H^{(L)}$ of the nodes in our arXiv citation network, which are then used for link prediction. Note that our purpose is to evaluate different scientific BERT models rather than GNNs, and we choose GraphSAGE due to its popularity. Different from earlier graph embedding methods that are transductive, GraphSAGE is an inductive graph representation learning algorithm and learns aggregation functions so that it can inductively generate embeddings of unseen nodes.
@@
$H^{(l)}=\sigma\left(\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}H^{(l-1)}W^{(l-1)}\right)\,,$ (2)
@@
$\begin{split}s(u,v)&=MLP\left(h_{u}^{(L)},h_{v}^{(L)}\right)\\ &=w_{2}\sigma\left(w_{1}\left[h_{u}^{(L)};h_{v}^{(L)}\right]+b_{1}\right)+b_{2}\,.\end{split}$ (3)
@@
$\begin{split}\mathcal{L}=-\sum_{u,v\in V}\Big{(}&y_{u\sim v}\log(s(u,v))+\\ &(1-y_{u\sim v})\log(1-s(u,v))\Big{)}\,,\end{split}$ (4)
@@
We consider the following 8 ways to obtain $H^{(0)}$.
  1. 1.
@@
  4. 4.
    The rest 5 feature matrices $H^{(0)}\in\mathbb{R}^{N\times 768}$ are encoded by each of the 5 scientific BERT models using the [CLS] tag pooling strategy on paper abstracts. For our models, we use AnnualBERT2020. By using different feature matrices encoded by the surveyed models, we can ascertain whether our AnnualBERT is the most useful encoder in predicting citation relationship between papers.
TABLE III: Experimental results for link prediction in the static arXiv citation network.
@@
GraphSAGE + AnnualBERT 0.991 (0.002) 0.955 (0.004) 0.967 (0.006) 0.942 (0.003) 0.955 (0.004)
V-A2 Experimental results
Table III reports the experimental results, providing a comprehensive comparison of various link prediction methods applied to our arXiv citation network. Firstly, the results indicate that some topological predictors—AA and RA—obtain the highest (possible) precision, highlighting their effectiveness in predicting connections between academic papers. However, these methods have rather low recall, suggesting that they miss many true positives. PPR appears to balance the best between precision and recall, considering especially its much higher recall than the other topological predictors.
@@ V-A2 Experimental results
Thirdly, using all the 5 different BERT models further improves AUC-ROC, implying that features encoded by them indeed provide additional information than subcategory does. Our AnnualBERT model stands out distinctly, achieving the highest score in AUC-ROC (0.991). Again, the same with the subcategory case, the increases in AUC-ROC stem from higher precision for all the BERT models except our AnnualBERT, as we can see that their recall remains constant around 0.91. Our AnnualBERT is the only model that improves both precision (0.967) and recall (0.942). These results demonstrate the effectiveness of our model in learning and encoding semantic information of papers among all the evaluated methods.
V-B Link prediction in temporal network
V-B1 Experimental setup

### v2

@@ 3.2. Pretraining AnnualBERT
The original BERT model is based on a bidirectional Transformer architecture and is pretrained through two objectives: Masked Language Model (MLM) and Next Sentence Prediction (NSP). In MLM, $15\%$ of the input tokens are masked, among which $80\%$ are replaced with the [MASK] token, $10\%$ are replaced with randomly selected tokens from the vocabulary, and the remaining $10\%$ remain unchanged. The objective is to predict the masked tokens using cross-entropy loss. The NSP task predicts whether two masked sentences follow one another in the original text. Building on BERT, the RoBERTa model (Liu et al., 2019) improves the pretraining by eliminating the NSP objective, which has been found less useful, and by implementing dynamic masking, where masked tokens change with each epoch, in contrast to the fixed masking used in BERT where masked tokens are fixed across epochs.
@@
With $\mathcal{M}_{2008}$, we perform, for each year $t>2008$, continual training of $\mathcal{M}_{t-1}$ on the corpus formed by papers published in $t$ for one epoch, resulting in $\mathcal{M}_{t}$. For instance, a continual training of $\mathcal{M}_{2008}$ on the corpus based on papers published in 2009 yields $\mathcal{M}_{2009}$, which is then used to develop $\mathcal{M}_{2010}$ using papers in 2010, in general following:
(1) $\mathcal{M}_{t}=\text{Continual-train}\left(\mathcal{M}_{t-1},\ \text{corpus}_{t}\right)\,.$
Through this approach, we allow each version of our AnnualBERT to capture evolving language and terminology specific to that year’s publications, so that we can represent the knowledge of a specific year in a more sophisticated way. By the time we extend the model to AnnualBERT2020, the corpora used for pretraining our AnnualBERT have grown to 2.7 billion words.
Finally, to further understand the effect of our pretraining method, we separately pretrain another RoBERTa model, denoted as $\overline{\mathcal{M}}$, on the entire, globally shuffled corpus (all papers published until 2020) for one epoch. This pretraining is completed in about 7 days on a server with 4 Nvidia L40 GPUs.
4. Performance on downstream NLP tasks
4.1. Task and dataset descriptions
@@
  1. (1)
    The SciERC dataset (Luan et al., 2018) annotates entities, relations, and coreference clusters in 500 abstracts from 12 AI conference and workshop proceedings indexed in the Semantic Scholar Corpus. It includes a total of $8,089$ distinct named entities.
  2. (2)
    The ScienceExam dataset (Smith et al., 2020) comprises entities from science exam questions aimed at students from 3rd to 9th grade, totaling $133,000$ entities.
  3. (3)
    The ScienceIE dataset (Augenstein et al., 2017) provides $9,946$ entities appeared in 500 scientific paragraphs.
@@
  1. (1)
    The SciERC dataset (Luan et al., 2018) includes $4,716$ annotated sentences categorized for classification according to relations such as USED-FOR, FEATURE-OF, PART-OF, etc.
  2. (2)
    The ACL-ARC dataset (Jurgens et al., 2018) comprises $1,941$ citation instances from 186 papers in the ACL Anthology Reference Corpus, with each instance annotated by domain experts with categories such as Future, Background, Motivation, etc.
  3. (3)
    The SciCite dataset (Cohan et al., 2019) encompasses $11,020$ citation instances from $6,627$ scholarly papers, with each citation classified into categories such as Background, Method, and Result Comparison.
@@
The last task is motivated by the increasing prevalence of interdisciplinary science (Gates et al., 2019) and its role in tackling important societal problems (Ledford, 2015).
@@
4.2. Experimental setup
We compare our AnnualBERT models with other scientific BERT models, including BioBERT (Lee et al., 2020), SciBERT (Beltagy et al., 2019), BioMedBERT (Gu et al., 2021), as well as the original BERT (Devlin et al., 2018). For our evaluations, we utilize the AnnualBERT model corresponding to the dataset’s year of release.
To ensure consistent comparisons across various BERT variants, we standardize the training configuration for each model. For the NER tasks, we adopt the NER model structure from SimpleTransformers, which incorporates a direct linear layer, and set the training parameters as follows: maximum input length of 512, training duration of 10 epochs, and batch size of 32. The best model parameters after each epoch are saved. For the CLS tasks, the vectors obtained using the [CLS] token pooling strategy are used as the input into a multilayer perceptron (MLP) for classification during the fine-tuning phase. The primary training parameters are set as follows: a training duration of 4 epochs, a batch size of 16, and weight decay at 0.01, and the best-performing model on the development set is saved every 100 steps. For the REL task, we use the dual tower structure (Reimers and Gurevych, 2019). We feed the abstracts of two arXiv papers into the identical BERT-like model and employ the [CLS] tag pooling strategy to obtain their respective feature vectors, $u$ and $v$, that effectively represent each abstract. We then concatenate $u$ and $v$ and their element-wise absolute difference $\left|u-v\right|$. This concatenated vector is then fed into a MLP layer for classification.
@@ 4.3. Experimental results
Table 2 presents the experimental results of various variants of BERT with two different methods of training our AnnualBERT: continual ($\mathcal{M}_{t}$) and “one-time” ($\overline{\mathcal{M}}$). Firstly, both variants of our AnnualBERT do not perform well on the NER tasks, and interestingly, neither do the other two models pretrained from scratch (SciBERT and BioMedBERT). This might be due to the tokenization strategy we use during the pretraining process. It may be also related to the fact that BioBERT, the best-performing model, is resulted from continual pretraining from the general Wikipedia and books corpus, which may have a better alignment with the benchmark datasets like the ScienceExam. Secondly, focusing still on the NER tasks, while both variants of AnnualBERT models have similar performances on the SciERC and ScienceExam datasets, the continual training version has a much better performance than the model from one-time training on the ScienceIE dataset.
Table 2. Experimental results of NLP tasks using different (scientific) BERT variants. Reported are F1 scores.
Task Field Dataset BERT BioBERT SciBERT BioMedBERT AnnualBERT $\mathcal{M}_{t}$ AnnualBERT $\overline{\mathcal{M}}$
CLS CS SciERC 85.96 87.18 87.21 86.96 87.60 86.49
CS ACL-ARC 76.22 75.97 80.32 79.53 79.88 81.80
Multi SciCite 85.20 86.63 86.33 86.33 85.77 85.09
Multi arXiv (major) 88.83 88.96 90.69 89.95 91.41 91.12
Multi arXiv (sub) 44.48 45.31 57.03 51.60 58.24 57.99
Cross-field identification arXiv 83.58 85.42 87.85 85.82 92.13 83.99
NER CS SciERC 59.69 63.61 62.55 61.84 60.96 60.80
Multi ScienceExam 80.53 82.88 82.13 80.69 77.69 78.28
Multi ScienceIE 33.56 34.86 33.97 34.40 33.95 30.93
REL Multi arXiv 86.79 87.64 89.18 87.44 89.52 90.43
@@
For the 3 sentence-level CLS tasks, on the other hand, our continually pretrained AnnualBERT has the best performance on the SciERC dataset and comparable performances to the other specialized BERT models on the ACL-ARC and SciCite datasets. In general, the performance differences among these models in the three tasks are rather small. However, significant performance gaps become evident for the abstract-level, three domain-specific tasks, where our continually pretrained AnnualBERT $\mathcal{M}_{t}$ performs the best. Particularly, for the sub-category prediction task, which involves classifying papers into one of the 172 sub-fields based on their abstracts, our AnnualBERT $\mathcal{M}_{t}$ significantly leads the other models including the one-time trained $\overline{\mathcal{M}}$. Notably, SciBERT also has a comparable performance in this task, likely due to its training corpus containing a substantial number of computer science papers. Finally, AnnualBERT (continual) has a dominant performance in the cross-field paper identification task, indicating its superior knowledge representation ability for interdisciplinary papers that stretch multiple domains, while the other models, despite fine-tuned, struggle with this task. Interestingly, while the AnnualBERT (one-time) shows relatively similar performances compared with its continual variant on most datasets, it performs significantly worse than the continual version and even other models on the task of cross-field paper identification. This disparity suggests that organizing the corpus by year, as done in the continual variant, may enhance the model’s ability to capture interdisciplinary knowledge. By incrementally exposing the model to cross-domain data over time, rather than feeding it all at once, our yearly-based continual training strategy appears to better equip language models for tasks spanning multiple domains. Taken together, these results substantiate the validity of our AnnualBERT models and the effectiveness of domain-specific pretraining for domain-specific tasks, in consistence with prior studies (Gu et al., 2021).
5. Citation link prediction
@@
Another perspective to assessing capabilities of our AnnualBERT models in learning and encoding semantic information of papers is through link prediction tasks in citation network. The rationale behind using this task lies in the nature of scientific discourse, where citations are formed by not only topical alignment but also nuanced interplay of numerous other factors like ideas and methodologies that are shared between two papers. Therefore, the ability of a language model in encoding abstracts into feature vectors that are highly predictive of whether two papers are related through citations reflects the model’s proficiency in grasping those complex semantic relationships, which in turn are instrumental for generating meaningful insights into domain-specific research (Peng et al., 2021a).
@@
As our setup requires a citation network between arXiv papers, we derive such a network from the Microsoft Academic Graph (MAG) dataset (Sinha et al., 2015). Doing so can avoid the otherwise extensive efforts of matching each referenced entry within the arXiv corpus, which may involve identifying the published version of a arXiv paper. To build our citation network, we first identify the corresponding MAG version of each arXiv paper, by extracting arXiv IDs from paper URLs provided by MAG, and then use the citation relationships between MAG papers to construct the network. Through this method, we identify nearly 1.5 million arXiv papers and 14.8 million citation relationships between them.
@@
5.1. Link prediction in static network
@@
5.1.1. Methods
Link prediction in networks has been extensively studied in the past (Liben-Nowell and Kleinberg, 2007; Ghasemian et al., 2020). Here we consider two families of approaches to predicting missing links: topological predictors and graph neural networks (GNN).
@@ 5.1.1. Methods
  4. (4)
    The Adamic/Adar (AA) index (Liben-Nowell and Kleinberg, 2007) defines the score as $s=\sum_{z\in\Gamma(u)\cap\Gamma(v)}\frac{1}{\log\left|\Gamma(z)\right|}$. It considers the importance of a common neighbor $z$, by giving more weight if $z$ has fewer connections, capturing the intuition that common neighbors in a less-populated field can be more significant in predicting a link between $u$ and $v$.
  5. (5)
    The resource allocation (RA) index (Zhou et al., 2009) defines the score as $s=\sum_{z\in\Gamma(u)\cap\Gamma(v)}\frac{1}{\left|\Gamma(z)\right|}$. It measures the likelihood of a link between two nodes based on the principle of resource allocation, where nodes share resources through their common neighbors.
  6. (6)
@@
One limitation of topological predictors is that they are unable to exploit node features, which in our context are the abstract of a paper. The rapid development of GNN has made it possible to integrate graph topology and node features for downstream tasks like link prediction. Here we leverage GraphSAGE (Hamilton et al., 2017) to generate embeddings $H^{(L)}$ of the nodes in our arXiv citation network, which are then used for link prediction. Note that our purpose is to evaluate different scientific BERT models rather than GNNs, and we choose GraphSAGE due to its popularity. Different from earlier graph embedding methods that are transductive, GraphSAGE is an inductive graph representation learning algorithm and learns aggregation functions so that it can inductively generate embeddings of unseen nodes.
@@
(2) $H^{(l)}=\sigma\left(\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}H^{(l-1)}W^{(l-1)}\right)\,,$
@@
(3) $\begin{split}s(u,v)&=MLP\left(h_{u}^{(L)},h_{v}^{(L)}\right)\\ &=w_{2}\sigma\left(w_{1}\left[h_{u}^{(L)};h_{v}^{(L)}\right]+b_{1}\right)+b_{2}\,.\end{split}$
@@
(4) $\mathcal{L}=-\sum_{u,v\in V}\left(y_{u\sim v}\log(s(u,v))+(1-y_{u\sim v})\log(1-s(u,v))\right)\,,$
@@
We consider the following 9 ways to obtain $H^{(0)}$.
  1. (1)
@@
  4. (4)
    The rest 6 feature matrices $H^{(0)}\in\mathbb{R}^{N\times 768}$ are encoded by each of the 6 scientific BERT models using the [CLS] tag pooling strategy on paper abstracts. For our models, we use AnnualBERT $\mathcal{M}_{2020}$ and $\overline{\mathcal{M}}$. By using different feature matrices encoded by the surveyed models, we can ascertain whether our AnnualBERT is the most useful encoder in predicting citation relationship between papers.
5.1.2. Experimental results
Table 3 reports the experimental results, providing a comprehensive comparison of various link prediction methods applied to our arXiv citation network. Firstly, the results indicate that some topological predictors—AA and RA—obtain the highest (possible) precision, highlighting their effectiveness in predicting connections between academic papers. However, these methods have rather low recall, suggesting that they miss many true positives. PPR appears to balance the best between precision and recall, considering especially its much higher recall than the other topological predictors.
Table 3. Experimental results for link prediction in the static arXiv citation network.
@@ 5.1.2. Experimental results
GraphSAGE+AnnualBERT $\mathcal{M}_{2020}$ 0.991 (0.002) 0.955 (0.004) 0.967 (0.006) 0.942 (0.003) 0.955 (0.004)
GraphSAGE+AnnualBERT $\overline{\mathcal{M}}$ 0.983 (0.002) 0.933 (0.003) 0.951 (0.005) 0.912 (0.003) 0.932 (0.003)
@@
Thirdly, using all the 5 different BERT models further improves AUC-ROC, implying that features encoded by them indeed provide additional information than subcategory does. Our AnnualBERT model $\mathcal{M}_{2020}$ stands out distinctly, achieving the highest score in AUC-ROC (0.991). Again, the same with the subcategory case, the increases in AUC-ROC stem from higher precision for all the BERT models except our AnnualBERT, as we can see that their recall remains constant around 0.91. Our AnnualBERT is the only model that improves both precision (0.967) and recall (0.942).
Fourthly, comparing one-time trained $\overline{\mathcal{M}}$ with continually trained $\mathcal{M}_{2020}$ reveals two findings: (1) The performance of $\overline{\mathcal{M}}$—despite pretrained on the in-domain corpus—is comparable with other models pretrained on out-of-domain corpora; and (2) our proposed continual pretraining based on temporally organized corpora is a more effective strategy, with the resulting model $\mathcal{M}_{2020}$ performing better than the one-time trained model $\overline{\mathcal{M}}$, even though they use identical pretraining corpus. In summary, these results demonstrate the effectiveness of our model in learning and encoding semantic information of papers among all the evaluated methods.
5.2. Link prediction in temporal network
5.2.1. Experimental setup